A beaker contains a total concentration of 0 62 molar of weak acetic acid with pka = 3.97 and pH = 2.22. Calculate the molar concentration of the undissociated (AH) form. Calculate the answer to three decimal places, do not specify units in the answer.

1
Expert's answer
2022-03-16T19:18:03-0400

( salt) =( 3.97×0.1)/2.22

8.8134/2 22

= 0.1788

And concentration of unneutralised Acetic acid is

( acid)=(0.62×0.1)/2.22

0.062/0.22= 0.0279)=

0.325 as final answer